OpenStudy (anonymous):

A train crossestwo person cyclying in the same direction, in 12 second and 15 seconds respectively.if the speed of the two cyclyst are 10 kmph and 12 kmph, then find the length and speed of train.

OpenStudy (anonymous):

dumbcow help me............

OpenStudy (anonymous):

speed = 20kmph andlength =100/3 meter but dont kno hw to solve

OpenStudy (dumbcow):

not sure on this one does it take 15 sec to pass both

OpenStudy (anonymous):

no 12 sec and 15 sec respectively

OpenStudy (dumbcow):

ok got it d = r*t r is speed of train d is distance of train t is time it take to pass each cyclist also need to convert from km/hr to m/s 1km/hr = 10/36 m/s cyclist 1: d = (r-10)(10/36)*(12) cyclist 2: d = (r-12)(10/36)*(15) to find r set 2 equations equal (r-10)(10/36)*12 = (r-12)(10/36)*(15) (10/36) cancels 12r -120 = 15r - 180 -3r = -60 r = 20 substitute this back in to find d d = (20-10)(10/36)*12 d = 1200/36 d = 100/3

OpenStudy (anonymous):

Let the speed of the train be v Thus, the relative speed of c1 = v-c1=v-10 the relative speed of c2 = v-c2=v-12 Now, speed = length of the train/time to cross the cycle v-10=L/12 and v-12=L/15 Therefore L/12+10 = L/15+12 L+120/12=L+180/15 15L+120*15 = 12L + 12*180 3L = 12*180 - 120*15 L = 4*180 - 40*15 =40(18-15) =40(3) 120.

OpenStudy (dumbcow):

you forgot to convert to m/s for speed because time is in sec

OpenStudy (anonymous):

I didn't

OpenStudy (anonymous):

Ya right, sorry!

OpenStudy (anonymous):

I don't know why they haven't given the option to delete answers!

OpenStudy (dumbcow):

amit did that help you understand

OpenStudy (anonymous):

yes